OpenStudy (anonymous):

Find the vertex of this parabola: y = -2x2 + 4x + 1 can someone help will give medal

OpenStudy (michele_laino):

the vertex V of a parabola has the subsequent coordinates: \[\Large V = \left( { - \frac{b}{{2a}}, - \frac{{{b^2} - 4ac}}{{2a}}} \right)\] where the coefficients a, b, and c are such that: y=ax^2+bx+c

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

please wait,I have made a typo, the rightformula is: \[\Large V = \left( { - \frac{b}{{2a}}, - \frac{{{b^2} - 4ac}}{{4a}}} \right)\]

OpenStudy (anonymous):

what do i do for the commas

OpenStudy (anonymous):

-8,8

OpenStudy (michele_laino):

so we have: \[\Large \begin{gathered} - \frac{b}{{2a}} = - \frac{4}{{2 \times \left( { - 2} \right)}} = ...? \hfill \\ \hfill \\ - \frac{{{b^2} - 4ac}}{{4a}} = - \frac{{16 - 4 \times \left( { - 2} \right) \times 1}}{{4 \times \left( { - 2} \right)}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

wait -1,-8

OpenStudy (michele_laino):

more explanation: the x-ccordinate of the vertex is: \[ - \frac{b}{{2a}} = - \frac{4}{{2 \times \left( { - 2} \right)}} = ...?\] whereas the y-coordinate of the vertex is: \[ - \frac{{{b^2} - 4ac}}{{4a}} = - \frac{{16 - 4 \times \left( { - 2} \right) \times 1}}{{4 \times \left( { - 2} \right)}} = ...?\] usually we separate those coordinates, by a comma, like this: \[V = \left( {x,y} \right)\]

OpenStudy (anonymous):

-1,-1

OpenStudy (michele_laino):

we have: \[ - \frac{b}{{2a}} = - \frac{4}{{2 \times \left( { - 2} \right)}} = \frac{4}{4} = 1\]

OpenStudy (michele_laino):

whereas: \[ - \frac{{{b^2} - 4ac}}{{4a}} = - \frac{{16 - 4 \times \left( { - 2} \right) \times 1}}{{4 \times \left( { - 2} \right)}} = - \frac{{24}}{{ - 8}} = 3.\]

OpenStudy (michele_laino):

so we can write: \[V = \left( {x,y} \right) = \left( {1,3} \right)\]

OpenStudy (anonymous):

ohhh ok i get thanks for helping me

OpenStudy (michele_laino):

thanks!